what is the weight of a block of ice 24in by 24in by 24in, if ice weighs 92 percent as much as water, and water weighs 62.5pounds per cubic feet?
Found 2 solutions by amoresroy, mananth:
Answer by amoresroy(333)

what is the weight of a block of ice 24in by 24in by 24in, if ice weighs 92 percent as much as water, and water weighs 62.5pounds per cubic feet?
24in = 2ft
Weight of block of ice = 2*2*2 *0.92*62.5
= 460 pounds

Answer by mananth(12270)

24 in = 2ft
..
volume of ice = 2*2*2 = 8 cu.ft
1cu.ft = 62.5 pounds of water
ice is= 92 % by weight of water
..
weight = 8 * 62.5 * 0.92
=460 pounds
...
m.ananth@hotmail.ca